OpenStudy (anonymous):

g(x)=6x^2-9x function domain and interval notation

OpenStudy (anonymous):

It's a polynomial so it's defined for all real numbers. Wonderful.

OpenStudy (anonymous):

You should look up interval notation to learn about it.

OpenStudy (anonymous):

In this case it is \[ (-\infty,\infty) \]

OpenStudy (anonymous):

I did I am just trying to figure out if I am to work it out and if so does it affect the infinity

OpenStudy (anonymous):

Well, for a polynomial, you know the natural domain is going to be all real numbers because there are no 'problem' areas. The main problem areas to look out for is when you could divide by zero or have a negative square root.

OpenStudy (anonymous):

For example \[ \frac{1}{x+1} \]Is not defined at \(-1\), so in interval notation we say:\[ (-\infty,-1)\cup (1,\infty) \]

OpenStudy (anonymous):

ok that makes more sense soif I had f(x)=13/(x+9) \[\infty, 9\]\[\infty, -9\]

OpenStudy (anonymous):

sorry I meant\[-\infty, -9) and (\infty, -9)\]

OpenStudy (anonymous):

If you wanna combine two intervals, generally we use \(\cup\), it does an OR operaion between two intervals, meaning any point in both intervals.

OpenStudy (anonymous):

any point in either interval^